OpenStudy (djjo):

Please help Algebra 1

OpenStudy (djjo):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4–x and y = 2x + 3 intersect are the solutions of the equation 4–x = 2x + 3. Part B: Make tables to find the solution to 4–x = 2x + 3. Take the integer values of x between −3 and 3.

OpenStudy (djjo):

I have part A I'm just confused with Part B

OpenStudy (djjo):

Is this right?: x y -3 64 = 2 -2 16 = -2 -1 4 = .25 0 0 = 8 1 -4 = 16 2 1/16 = 32 3 1/64 = 64

OpenStudy (anonymous):

I believe what B is saying is make tables and it has to range between -3 and 3

OpenStudy (anonymous):

Make 3 columns on your table: x, 4-x, and 2x+3

OpenStudy (anonymous):

Use x value from -3 to 3, and look to where 4-x is equal to 2x+3. That value of x will be your solution.

OpenStudy (djjo):

ok thx

OpenStudy (anonymous):

Eg. x 4-x 2x+3 -3 7 -3 and so on

OpenStudy (djjo):

thx so much @ospreytriple

OpenStudy (anonymous):

You're welcome